OpenStudy (swissgirl):

I need an example of a union of 2 subgroups that isnt a subgroup

OpenStudy (swissgirl):

It sounds more confusing than it is Basically H and K are subgroups of G but \( H \cup K \) isnt a subgroup of G

OpenStudy (anonymous):

what groups do you know?

OpenStudy (swissgirl):

I know nothing

OpenStudy (swissgirl):

Cyclic groups

OpenStudy (swissgirl):

and that is it

OpenStudy (anonymous):

if you pick two subgroups of \(S_3\) and take there union it will not be a subgroup

OpenStudy (anonymous):

but that one is not cyclic how about the group \(\mathbb{Z}\) under addition

OpenStudy (swissgirl):

ohhhhh It doesnt have to be cyclic

OpenStudy (anonymous):

subgroups of \(\{\mathbb{Z},+\} \) look like what?

OpenStudy (anonymous):

Yeah, I can't think how this is even possible...

OpenStudy (swissgirl):

it is. I am in middle of working on \( (\mathbb{Z_4}, +) \)

OpenStudy (anonymous):

yeah. I am really weak on Set Theory. Next thing to brush up on...

OpenStudy (swissgirl):

I am weak on everything -_-

OpenStudy (anonymous):

subgroups of \(\{\mathbb{Z},+\}\) look like \(n\mathbb{Z}\)

OpenStudy (anonymous):

\[ A= 5 Z\\ B=7 Z\\ 5 \in A \cup B\\ 7 \in A \cup B\\ but\\ 5+7 \notin A \cup B\\ \]

OpenStudy (swissgirl):

huh I had a totally diff example

OpenStudy (swissgirl):

I dont get it what u guys r doing

OpenStudy (anonymous):

so take for example the union of two such subgroups, say \(7\mathbb{Z}\) and \(5\mathbb{Z}\)

OpenStudy (anonymous):

oh wow i am being a copy cat, i guess it was the easiest answer

OpenStudy (swissgirl):

lol

OpenStudy (anonymous):

is it clear what i wrote (after @eliassaab wrote the same thing first?)

OpenStudy (swissgirl):

Like i dont get ur notation. They dont show any of that in my book

OpenStudy (swissgirl):

But I have another exxample

OpenStudy (anonymous):

as subgroup of \(\mathbb{Z}\) under addition looks like all multiples of some number say 5 so it would be the set \(\{...-15,-10,-5,0,5,10,15,...\}\)

OpenStudy (swissgirl):

okk

OpenStudy (anonymous):

you can check that this is a subgroup easily now pick another such subgroup because apparently i have no originality pick \[7\mathbb{Z}=\{...-21,-14,-7,0,7,14,21,...\}\]

OpenStudy (anonymous):

their union is what? all multiples of 5 or 7 but it is not even closed, since \(5+7\) is not in there

OpenStudy (swissgirl):

ohhhhh i get ittttttt

OpenStudy (swissgirl):

Thanksss idk all i see in my book are these mod examples so i just used one of those

OpenStudy (swissgirl):

Thankkkssssss guys